OpenStudy (anonymous):

find the domain of the logarithmic function: f(x)=ln(x^2+3x-4)

OpenStudy (anonymous):

basically, whats inside the parentheses cannot be equal to or less than zero

OpenStudy (anonymous):

or put an equivalent way, solve \(x^2+3x-4>0\)

OpenStudy (anonymous):

so i factor the problem, and get -4 and 1, so is that the domain? or can it not be negative...

OpenStudy (anonymous):

\(x^2+3x-4\) is a parabola that opens up. you found the zeros via \((x+4)(x-1)=0\) so your parabola will be negative between the zeros and positive outside them

OpenStudy (anonymous):



OpenStudy (anonymous):

you have to make sure it is positive, so your answer should look like \(x<-1\) or \(x>4\) or perhaps \((-\infty, -1)\cup (4,\infty)\)

OpenStudy (anonymous):

thanks! that's exactly what i was looking for

OpenStudy (anonymous):

Just a little typo: x < -4 or x > 1

OpenStudy (anonymous):

i'll say it was a typo. comletey wrong in fact. cholorphll has the right answer

OpenStudy (anonymous):

\[(-\infty, -4)\cup (1,\infty)\]

OpenStudy (anonymous):

@satellite73 I just jump in at the last minute to switch the sign!